Write a balanced chemical equation for the reaction described in question 2, assuming that the only products formed in the reaction are solid silver metal, carbon dioxide gas, and water vapor. Include the phases of each reactant and product.

Question 2. (A sample of solid silver oxide with a mass of 0.114 grams was reduced to elemental silver by heating under a flow of methane gas, CH4, Determine the mass percent of each element in silver oxide and the empirical formula of silver oxide if the reaction produced 0.106grams of silver.)
Answer #1

Q.2

Mass of silver present in the oxide = 0.106 g.

Mass of oxygen present = ( 0.114 - 0.106) g = 0.008 g.

Percentage of silver =

=

= 92.98 %.

Percentage of oxygen

= ​​​​​​

=

= 7.02 % .

Now,

Mole ratio of Ag and O

= Ag : O

= :

= :

= 0.00098 : 0.0005

= (0.00098/0.0005) : (0.0005/0.0005)

= 1.96 : 1

2:1

Hence, empirical formula is Ag2O.

Balanced equation is

4 Ag2O (s) + CH4 (g) 8 Ag (s) + CO2 (g) + 2 H2O (g)